# TWSSenti: A Novel Hybrid Framework for Topic-Wise Sentiment Analysis on Social Media Using Transformer Models

### v1

@@ Abstract
Sentiment analysis is a crucial task in natural language processing (NLP) that enables the extraction of meaningful insights from textual data, particularly from dynamic platforms like Twitter and IMDB. This study explores a hybrid framework combining transformer-based models, specifically BERT, GPT-2, RoBERTa, XLNet, and DistilBERT, to improve sentiment classification accuracy and robustness. The framework addresses challenges such as noisy data, contextual ambiguity, and generalization across diverse datasets by leveraging the unique strengths of these models. BERT captures bidirectional context, GPT-2 enhances generative capabilities, RoBERTa optimizes contextual understanding with larger corpora and dynamic masking, XLNet models dependency through permutation-based learning, and DistilBERT offers efficiency with reduced computational overhead while maintaining high accuracy. We demonstrate text cleaning, tokenization, and feature extraction using Term Frequency Inverse Document Frequency (TF-IDF) and Bag of Words (BoW), ensure high-quality input data for the models. The hybrid approach was evaluated on benchmark datasets Sentiment140 and IMDB, achieving superior accuracy rates of 94% and 95%, respectively, outperforming standalone models. The results validate the effectiveness of combining multiple transformer models in ensemble-like setups to address the limitations of individual architectures. This research highlights its applicability to real-world tasks such as social media monitoring, customer sentiment analysis, and public opinion tracking which offers a pathway for future advancements in hybrid NLP frameworks.
@@ 1 Introduction
Traditional sentiment analysis approaches, including machine learning algorithms like Support Vector Machines (SVM) and Naïve Bayes, as well as lexicon-based methods, have been widely employed for text classification [36, 25, 32]. However, these methods often struggle to handle the complexity of natural language, such as contextual nuances, ambiguous expressions, and informal language commonly used on platforms like Twitter [12, 18]. With the advent of deep learning and, more recently, transformer-based models, the landscape of sentiment analysis has undergone a significant transformation. Despite their success, these models are not without limitations. Challenges such as overfitting, computational inefficiency, and poor generalization across datasets persist, particularly when applied to real-world, large-scale datasets like Sentiment140 and IMDB. The motivation behind this research stems from the need to address these limitations and further enhance the capabilities of transformer-based models for sentiment analysis. Individual transformer models, while powerful, often fail to capture all the intricacies of textual data due to their reliance on specific training paradigms. Additionally, standalone models are prone to overfitting when trained on smaller or imbalanced datasets that lead to suboptimal performance in real-world applications.
@@ 2.1 Traditional Machine Learning Approaches
Traditional machine learning techniques were among the earliest methods employed for sentiment analysis. Algorithms like Support Vector Machines (SVM), Naive Bayes, and decision trees were widely used for text classification tasks [20, 44, 34]. These approaches relied heavily on handcrafted features, such as n-grams, Bag of Words (BoW), and Term Frequency-Inverse Document Frequency (TF-IDF), to represent textual data numerically [22, 23]. For instance, BoW represented text as a sparse vector of word occurrences, while TF-IDF weighted words based on their importance across a document corpus [21]. Although these methods provided foundational insights into text classification, they were inherently limited by their inability to capture semantic relationships and contextual dependencies in text. For example, while SVMs and Naive Bayes achieved reasonable accuracy, their performance declined significantly on larger or more complex datasets. The inability to handle long-range dependencies in text further constrained their scalability. Moreover, these methods required extensive feature engineering, making them labor-intensive and less adaptive to changes in data or domain. The authors [7, 30] highlighted that traditional machine learning approaches often struggled to generalize well across diverse datasets, thereby limiting their applicability to real-world sentiment analysis tasks. These limitations underscored the need for more advanced, context-aware models capable of learning features automatically.
@@ 2.3 Transformer-Based Models
The advent of transformer-based architectures revolutionized sentiment analysis by introducing mechanisms that captured long-range dependencies and bidirectional context. Models like BERT, GPT-2, RoBERTa, XLNet, and DistilBERT set new benchmarks in text classification tasks [29, 37]. BERT, for instance, utilized a bidirectional transformer to capture context from both directions, enabling a deeper understanding of sentence structure and semantics [45]. RoBERTa improved upon BERT by training on larger datasets and employing dynamic masking techniques, resulting in enhanced performance on sentiment analysis tasks [19]. GPT-2, with its autoregressive structure, excelled in generative tasks but was less effective for classification tasks due to its unidirectional context modeling [26]. XLNet introduced permutation-based modeling, which enhanced bidirectional learning by considering all possible word orderings in a sentence [4]. This study made XLNet particularly robust for long-form text, such as IMDB movie reviews, where understanding global context was critical. The authors demonstrated that transformer-based models like BERT and RoBERTa consistently outperformed traditional and deep learning models, achieving state-of-the-art results on datasets like Sentiment140 and IMDB [39]. However, transformer models are not without limitations. These architectures are computationally intensive, requiring significant resources for training and fine-tuning. They are also prone to overfitting on small or imbalanced datasets, where a lack of data diversity can limit generalization. Moreover, individual transformer models often struggle with domain-specific challenges, such as sarcasm, slang, or abbreviations common in social media text. While their self-attention mechanisms allow for expressive encoding of relationships within text, they lack adaptability to diverse linguistic and cultural contexts without domain-specific fine-tuning.
@@ 3.8 Proposed Algorithm
0:  Accuracy $A$ and Confusion Matrix $CM$. Initialization:
@@ 3.9 Initialization
$D.\text{text}$ Textual data column in dataset $D$.
$D.\text{sentiment}$ Sentiment labels corresponding to textual data in dataset $D$.
@@
PretrainModel Pretrained transformer model used for sequence classification.
@@
$A$ Accuracy of the model, calculated as the ratio of correct predictions to total predictions.
$CM$ Confusion matrix, showing the distribution of true positives, false positives, false negatives, and true negatives.
@@ 4.2.2 Impact of Preprocessing and Feature Engineering
A critical factor in the success of the hybrid model was the comprehensive preprocessing pipeline, which ensured high-quality input data. Text cleaning removed irrelevant components such as URLs, mentions, and special characters, reducing noise in the datasets. Tokenization, stopword removal, and stemming or lemmatization standardized the input, enabling better feature extraction. The use of TF-IDF as the primary feature extraction technique played a pivotal role in quantifying the importance of terms, particularly in distinguishing between sentiment classes. This step ensured that the model focused on meaningful textual patterns, significantly improving classification performance. These insights highlight the importance of preprocessing and feature engineering as foundational steps in achieving high accuracy and robustness in sentiment analysis.
@@ 4.2.4 Theoretical Implications
4.3 Enhancements to Computational Efficiency and Deployment Feasibility
@@
Model Training Time
per Epoch (min)
Inference Time
(ms/batch)
GPU Memory
Usage (GB)
@@
4.4 Comparative Analysis of Computational Costs and Model Performance
@@
Approach Accuracy
Accuracy
Inference
Time (ms)
Computational
Cost
@@ 4.7 Comparison to the State-of-the-Art Methods
Our proposed hybrid approach, which integrates BERT, GPT-2, RoBERTa, XLNet, and DistilBERT, achieves superior performance with an accuracy of $95.00\%$ on IMDB and $94.00\%$ on Sentiment140. This surpasses the best-performing alternative model, RoBERTa-GRU by  [39], which achieves $94.63\%$ on IMDB ($+0.37\%$ improvement) and $89.59\%$ on Sentiment140 ($+4.41\%$ improvement). The significant increase in accuracy on Sentiment140 demonstrates the robustness of our model in handling noisy social media text. Similarly,  [28] proposed a RoBERTa-BiLSTM model, which achieved $92.36\%$ on IMDB ($+2.64\%$ improvement) and $82.25\%$ on Sentiment140 ($+11.75\%$ improvement). While this model provides a competitive performance, its lower accuracy on Sentiment140 suggests limitations in generalizing across diverse textual sources.
[8] experimented with multiple hybrid transformer models, including DistilBERT-BiLSTM, BERT-BiLSTM, and RoBERTa-CNN-BiLSTM, achieving maximum testing accuracies of $81.00\%$, $79.00\%$, and $77.00\%$ on Sentiment140. Our model significantly outperforms these approaches, with an improvement of up to $+17.00\%$ over DistilBERT-BiLSTM, indicating that a deeper fusion of multiple transformers provides enhanced feature extraction and sentiment classification.  [31] introduced a semi-supervised boosting framework (SSSA), obtaining $79.30\%$ on IMDB and $80.30\%$ on Sentiment140. While their approach benefits from semi-supervised learning, its performance remains substantially lower than our hybrid method, with a $+15.70\%$ increase on IMDB and a $+13.70\%$ increase on Sentiment140. This highlights the advantage of using fully supervised transformer ensembles for large-scale sentiment classification.
@@
Ref. Model Dataset Testing
Accuracy (%)

### v2

# arXiv:2504.09896v2[cs.CL] 26 Feb 2026

@@ Abstract
Sentiment analysis is a crucial task in natural language processing (NLP) that enables the extraction of meaningful insights from textual data, particularly from dynamic platforms like Twitter and IMDB. This study explores a hybrid framework combining transformer-based models, specifically BERT, GPT-2, RoBERTa, XLNet, and DistilBERT, to improve sentiment classification accuracy and robustness. The framework addresses challenges such as noisy data, contextual ambiguity, and generalization across diverse datasets by leveraging the unique strengths of these models. BERT captures bidirectional context, GPT-2 enhances generative capabilities, RoBERTa optimizes contextual understanding with larger corpora and dynamic masking, XLNet models dependency through permutation-based learning, and DistilBERT offers efficiency with reduced computational overhead while maintaining high accuracy. We demonstrate text cleaning, tokenization, and feature extraction using Term Frequency Inverse Document Frequency (TF-IDF) and Bag of Words (BoW), ensuring high-quality input data for the models. The hybrid approach was evaluated on benchmark datasets Sentiment140 and IMDB, achieving superior accuracy rates of 94% and 95%, respectively, outperforming standalone models. Experimental results show that TWSSenti achieves accuracies of 94% on Sentiment140 and 95% on IMDB. This corresponds to a relative improvement of 3.2% and 2.7%, respectively, over strong transformer baselines such as BERT and XLNet, and more than a 10% gain compared with traditional machine learning models. These improvements demonstrate the effectiveness of the proposed hybrid fusion approach for fine-grained sentiment analysis. The results validate the effectiveness of combining multiple transformer models in ensemble-like setups to address the limitations of individual architectures. This research highlights its applicability to real-world tasks such as social media monitoring, customer sentiment analysis, and public opinion tracking which offer a pathway for future advancements in hybrid NLP frameworks.
@@ 1 Introduction
Traditional sentiment analysis approaches, including machine learning algorithms like Support Vector Machines (SVM) and Naïve Bayes, as well as lexicon-based methods, have been widely employed for text classification [38, 27, 45, 34]. However, these methods often struggle to handle the complexity of natural language, such as contextual nuances, ambiguous expressions, and informal language commonly used on platforms like Twitter [13, 19, 44]. With the advent of deep learning and, more recently, transformer-based models, the landscape of sentiment analysis has undergone a significant transformation. Despite their success, these models are not without limitations. Challenges such as overfitting, computational inefficiency, and poor generalization across datasets persist, particularly when applied to real-world, large-scale datasets like Sentiment140 and IMDB. The motivation behind this research stems from the need to address these limitations and further enhance the capabilities of transformer-based models for sentiment analysis. Individual transformer models, while powerful, often fail to capture all the intricacies of textual data due to their reliance on specific training paradigms. Additionally, standalone models are prone to overfitting when trained on smaller or imbalanced datasets that lead to suboptimal performance in real-world applications.
@@ 2.1 Traditional Machine Learning Approaches
Traditional machine learning techniques were among the earliest methods employed for sentiment analysis. Algorithms like Support Vector Machines (SVM), Naive Bayes, and decision trees were widely used for text classification tasks [21, 52, 36, 43]. These approaches relied heavily on handcrafted features, such as n-grams, Bag of Words (BoW), and Term Frequency-Inverse Document Frequency (TF-IDF), to represent textual data numerically [23, 24]. For instance, BoW represented text as a sparse vector of word occurrences, while TF-IDF weighted words based on their importance across a document corpus [22]. Although these methods provided foundational insights into text classification, they were inherently limited by their inability to capture semantic relationships and contextual dependencies in text. For example, while SVMs and Naive Bayes achieved reasonable accuracy, their performance declined significantly on larger or more complex datasets. The inability to handle long-range dependencies in text further constrained their scalability. Moreover, these methods required extensive feature engineering, making them labor-intensive and less adaptive to changes in data or domain. The authors [8, 32] highlighted that traditional machine learning approaches often struggled to generalize well across diverse datasets, thereby limiting their applicability to real-world sentiment analysis tasks. These limitations underscored the need for more advanced, context-aware models capable of learning features automatically.
@@ 2.3 Transformer-Based Models
The advent of transformer-based architectures revolutionized sentiment analysis by introducing mechanisms that captured long-range dependencies and bidirectional context. Models like BERT, GPT-2, RoBERTa, XLNet, and DistilBERT set new benchmarks in text classification tasks [31, 39]. BERT, for instance, utilized a bidirectional transformer to capture context from both directions, enabling a deeper understanding of sentence structure and semantics [53]. RoBERTa improved upon BERT by training on larger datasets and employing dynamic masking techniques, resulting in enhanced performance on sentiment analysis tasks [20]. GPT-2, with its autoregressive structure, excelled in generative tasks but was less effective for classification tasks due to its unidirectional context modeling [28, 43, 46]. XLNet introduced permutation-based modeling, which enhanced bidirectional learning by considering all possible word orderings in a sentence [5, 45]. This study made XLNet particularly robust for long-form text, such as IMDB movie reviews, where understanding global context was critical. The authors demonstrated that transformer-based models like BERT and RoBERTa consistently outperformed traditional and deep learning models, achieving state-of-the-art results on datasets like Sentiment140 and IMDB [41]. However, transformer models are not without limitations. These architectures are computationally intensive, requiring significant resources for training and fine-tuning. They are also prone to overfitting on small or imbalanced datasets, where a lack of data diversity can limit generalization. Moreover, individual transformer models often struggle with domain-specific challenges, such as sarcasm, slang, or abbreviations common in social media text. While their self-attention mechanisms allow for expressive encoding of relationships within text, they lack adaptability to diverse linguistic and cultural contexts without domain-specific fine-tuning.
@@ 3.8 Proposed Algorithm
0: Accuracy AA and Confusion Matrix CMCM. Initialization:
@@ 3.9 Initialization
$D.\texttt{text}$ Textual data column in dataset $D$.
$D.\texttt{sentiment}$ Sentiment labels corresponding to textual data in dataset $D$.
@@
PretrainedModel Pretrained transformer model used for sequence classification.
@@
AA Accuracy of the model, calculated as the ratio of correct predictions to total predictions.
CMCM Confusion matrix, showing the distribution of true positives, false positives, false negatives, and true negatives.
@@ 4.2.2 Impact of Preprocessing and Feature Engineering
A critical factor in the success of the hybrid model was the comprehensive preprocessing pipeline, which ensured high-quality input data. Text cleaning removed irrelevant components such as URLs, mentions, and special characters, reducing noise in the datasets. Tokenization, stopword removal, and stemming or lemmatization standardized the input, enabling better feature extraction. The use of TF-IDF as the primary feature extraction technique played a pivotal role in quantifying the importance of terms, particularly in distinguishing between sentiment classes. This step ensured that the model focused on meaningful textual patterns, significantly improving classification performance. These insights highlight the importance of preprocessing and feature engineering as foundational steps in achieving high accuracy and robustness in sentiment analysis. Although the experimental evaluation is conducted using two commonly used datasets—Sentiment140 and IMDB—these datasets capture two distinct and complementary language styles. Sentiment140 represents short, informal, noisy user-generated content commonly found on social media platforms, whereas IMDB reviews consist of longer, more structured sentences with richer contextual dependencies. Evaluating TWSSenti on both datasets therefore demonstrates its ability to generalize across different input lengths, writing styles, and sentiment expressions. Furthermore, because the proposed fusion mechanism operates at the probability level and does not rely on domain-specific lexicons or handcrafted rules, it can be readily applied to new domains such as finance, healthcare, customer service logs, product reviews, and political discourse. Future extensions may incorporate domain adaptation or continual pre-training to further validate the framework across specialized datasets.
4.2.3 Interpretability and Explainability of the Hybrid Model
Although deep transformer architectures are often viewed as black-box models, the hybrid configuration of TWSSenti affords several advantages for interpretability and decision transparency in sentiment analysis. Each transformer within the ensemble provides token-level attention scores that highlight the most influential words or phrases contributing to the final sentiment classification. By aggregating these attention maps across models such as BERT, RoBERTa, XLNet, and GPT-2, the hybrid system produces more stable and consistent explanations compared to relying on a single architecture.
Furthermore, the soft-probability fusion mechanism offers an additional layer of interpretability by revealing cross-model agreement. When multiple transformers assign high probability to the same sentiment class, the decision is more reliable and easier to justify. Conversely, disagreements among models can flag potentially ambiguous or context-sensitive inputs, enabling users to inspect difficult cases more thoroughly.
The ensemble design also reduces bias inherent in individual models, as each transformer captures different linguistic patterns—BERT excels at bidirectional context, GPT-2 highlights generative contextual flows, and XLNet captures long-range dependencies. Their combined perspectives lead to a more comprehensive understanding of why the system assigns a particular sentiment label. Together, these mechanisms improve transparency and support more informed decision-making in applications such as customer feedback analysis, social media monitoring, and opinion mining.
4.2.4 Optimization Strategies for Practical Deployment
Although transformer-based models offer strong predictive performance, training and maintaining multiple large architectures can be computationally expensive. To improve practicality in real-world settings, several optimization techniques can be applied to the TWSSenti framework. First, model pruning can be used to remove redundant attention heads or feed-forward connections, thereby reducing the computational burden while retaining most of the predictive accuracy. Second, quantization techniques—such as converting weights and activations from 32-bit floating point to 8-bit integer representations—significantly lower memory requirements and accelerate inference on edge devices. Third, knowledge distillation enables the transfer of knowledge from a set of large “teacher” models to a single lightweight “student” model, offering a substantial reduction in model size with minimal loss in performance. By incorporating these optimization strategies, the hybrid framework can be deployed more efficiently in latency-sensitive or resource-constrained environments without sacrificing its overall sentiment classification accuracy.
@@ 4.2.6 Theoretical Implications
4.2.7 Enhancements to Computational Efficiency and Deployment Feasibility
@@
Model Training Time per Epoch (min) Inference Time (ms/batch) GPU Memory Usage (GB)
@@
4.2.8 Comparative Analysis of Computational Costs and Model Performance
@@
Approach Accuracy (IMDB) Accuracy (Sentiment140) Inference Time (ms) Computational Cost
@@ 4.5 Comparison to the State-of-the-Art Methods
Our proposed hybrid approach, which integrates BERT, GPT-2, RoBERTa, XLNet, and DistilBERT, achieves superior performance with an accuracy of $95.00\%$ on IMDB and $94.00\%$ on Sentiment140. This surpasses the best-performing alternative model, RoBERTa-GRU by  [41], which achieves $94.63\%$ on IMDB ($0.37\%$ improvement) and $89.59\%$ on Sentiment140 ($4.41\%$ improvement). The significant increase in accuracy on Sentiment140 demonstrates the robustness of our model in handling noisy social media text. Similarly,  [30] proposed a RoBERTa-BiLSTM model, which achieved $92.36\%$ on IMDB ($2.64\%$ improvement) and $82.25\%$ on Sentiment140 ($11.75\%$ improvement). While this model provides a competitive performance, its lower accuracy on Sentiment140 suggests limitations in generalizing across diverse textual sources.
[9] experimented with multiple hybrid transformer models, including DistilBERT-BiLSTM, BERT-BiLSTM, and RoBERTa-CNN-BiLSTM, achieving maximum testing accuracies of $81.00\%$, $79.00\%$, and $77.00\%$ on Sentiment140. Our model significantly outperforms these approaches, with an improvement of up to $17.00\%$ over DistilBERT-BiLSTM, indicating that a deeper fusion of multiple transformers provides enhanced feature extraction and sentiment classification.  [33] introduced a semi-supervised boosting framework (SSSA), obtaining $79.30\%$ on IMDB and $80.30\%$ on Sentiment140. While their approach benefits from semi-supervised learning, its performance remains substantially lower than our hybrid method, with a $15.70\%$ increase on IMDB and a $13.70\%$ increase on Sentiment140. This highlights the advantage of using fully supervised transformer ensembles for large-scale sentiment classification.
@@
Ref. Model Dataset Testing Accuracy (%)